OpenStudy (anonymous):

An HP printer prints 60 copies per minute. However, it takes 10 minutes to heat up before it prints any copies. Assume the copier runs continuously and steadily. 1) Write an equation for the number of copies, n(t), as a function of the time, t, in slope-intercept form, using function notation. 2) How long does it take to print 900 copies? 3) How many copies will the printer make in 90 mins?

OpenStudy (anonymous):

n(t)=60*(t-10) put n(t)=900 gives t=25 60*(90-10)=80*60=4800

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."